# Experience of Training a 1.7B-Parameter LLaMa Model From Scratch

### v2

@@ Abstract
Pretraining large language models is a complex endeavor influenced by multiple factors, including model architecture, data quality, training continuity, and hardware constraints. In this paper, we share insights gained from the experience of training DMaS-LLaMa-Lite, a fully open source, 1.7-billion-parameter, LLaMa-based model, on approximately 20 billion tokens of carefully curated data. We chronicle the full training trajectory, documenting how evolving validation loss levels and downstream benchmarks reflect transitions from incoherent text to fluent, contextually grounded output. Beyond pretraining, we extend our analysis to include a post-training phase focused on instruction tuning, where the model was refined to produce more contextually appropriate, user-aligned responses. We highlight practical considerations such as the importance of restoring optimizer states when resuming from checkpoints, and the impact of hardware changes on training stability and throughput. While qualitative evaluation provides an intuitive understanding of model improvements, our analysis extends to various performance benchmarks, demonstrating how high-quality data and thoughtful scaling enable competitive results with significantly fewer training tokens. By detailing these experiences and offering training logs, checkpoints, and sample outputs, we aim to guide future researchers and practitioners in refining their pretraining strategies. The training script is available on Github at https://github.com/McGill-DMaS/DMaS-LLaMa-Lite-Training-Code. The model checkpoints are available on Huggingface at https://huggingface.co/collections/McGill-DMaS/dmas-llama-lite-6761d97ba903f82341954ceb.
I Introduction
Large language models (LLMs) have achieved unprecedented performance across a variety of natural language understanding and generation tasks. These advancements have been driven by innovations in model architectures, pretraining methodologies, and the availability of large-scale, high-quality data [1, 2, 3]. However, despite significant progress, many aspects of LLM training remain underexplored. Practical challenges—ranging from data curation and training stability to qualitative and quantitative evaluation—play a critical role in determining model performance but are often overlooked in favor of final benchmark results.
In this paper, we present our experience pretraining and post-training DMaS-LLaMa-Lite, a 1.7-billion-parameter LLaMa-based model, on 20 billion tokens of carefully curated data. Unlike many existing studies that focus solely on the final outputs of large-scale training, we emphasize the training process itself and the insights gleaned from it. Specifically, we document the following key aspects of our work:
@@ I Introduction
    Post-Training Instruction Tuning We refine the pretrained model using parameter-efficient fintuning, leveraging high-quality instruction datasets. We documented the model’s behavioral change from the process and the model’s sensitivity to the prompt.
Our findings collectively offer insights into the practical challenges and trade-offs encountered during the pretraining of LLMs. By sharing detailed training logs, checkpoint artifacts, and sample completions, we aim to support reproducibility and provide guidance for both researchers and practitioners seeking to optimize their own training strategies. Importantly, this work serves as a foundation for future studies that may uncover additional best practices, methodological refinements, or novel evaluation techniques in the pretraining of large language models.
The remainder of this paper is organized as follows: Section  II describes the model architecture, training setup, and data preprocessing. Section  III presents our findings from the pretraining experience, including quantitative and qualitative evaluations, and a discussion of broader implications. Section  IV details the post-training phase, including instruction tuning and qualitative analysis. Finally, Section  V concludes with key takeaways and directions for future work.
II Model Configuration and Training Setup
II-A Model Architecture
Our model, DMaS-LLaMa-Lite, is a 1.7B-parameter (1716M) variant of LLaMa [3] adapted to use a GPT-2 tokenizer. The configuration details are presented in Table I. The model features 36 transformer layers, each with a hidden size of 2048 and an intermediate size of 5120 in the feed-forward layers. We employ 32 attention heads, grouped into 8 key-value heads, and use a rotary positional embedding scheme. Norm layers use RMS normalization [4], and the model’s non-linear activation is SiLU [5]. The model is trained on tokens produced by the GPT-2 tokenizer.
@@ II-A Model Architecture
TABLE I: Model configuration for the 1.7B-parameter DMaS-LLaMa-Lite model.
II-B Training Data and Preprocessing
The training was conducted using a subset of the HuggingFaceFW/fineweb-edu 100BT [6] dataset, derived from the broader FineWeb dataset. FineWeb-Edu comprises approximately 1.3 trillion tokens of high-quality educational content, curated through a meticulous multi-stage filtering process. FineWeb itself is a 15-trillion-token dataset sourced from 96 Common Crawl snapshots, designed to improve LLM performance through advanced filtering and deduplication methods.
FineWeb-Edu emphasizes educational text, selected using a classifier trained on synthetic annotations generated by Llama-3-70B-Instruct [7]. This classifier was fine-tuned to prioritize grade-school and middle-school-level knowledge while excluding overly technical content. The dataset employs custom filters targeting document structure and content quality, ensuring a high proportion of semantically coherent and educationally relevant text.
@@ II-B Training Data and Preprocessing
II-C Optimizer and Learning Rate Schedule
@@
TABLE II: Evaluation Scores for Different Training Steps (Updated)
@@
II-D Training Procedure and Checkpoints
The training was conducted on one or two RTX A6000 GPUs, reflecting realistic constraints in research and production environments. When using two GPUs, PyTorch’s DistributedDataParallel (DDP) [9] was employed to ensure efficient data and model parallelism. Transitions in the hardware setup (e.g., moving from a single RTX A6000 GPU to dual RTX A6000 GPUs) served as natural breakpoints, allowing us to examine how changes in training continuity, combined with optimizer state restoration or omission, affected validation loss and model quality.
III Pretraining
III-A Validation Loss Trajectories and Training Metrics
III-B Qualitative Evaluation of Generated Text
TABLE III: Generated completions at different training steps.
Checkpoint Generated Text
Prompt ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note”
2700 ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the impressive size and precision of the structures, marveling at how such feats of engineering were accomplished thousands of years ago without modern tools.”
3300 ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the mystery surrounding their construction, as well as the stunning alignment with celestial bodies.”
5100 ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the grandeur and historical significance of these monuments, which continue to captivate people from around the globe.”
15700 ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note their intricate design, which reflects the advanced knowledge of ancient Egyptian engineers and architects.”
35000 ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the awe-inspiring scale and the enduring mystery of how they were constructed with such precision.”
43250 ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the breathtaking beauty and the remarkable testament to the ingenuity of ancient civilizations.”
Prompt ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for”
2700 ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its enigmatic smile and the use of sfumato technique to create a lifelike expression.”
3300 ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its detailed background, which demonstrates da Vinci’s mastery of perspective.”
5100 ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its captivating expression, which has intrigued viewers and art historians for centuries.”
15700 ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for the subtle play of light and shadow on the subject’s face, a hallmark of Renaissance art.”
35000 ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its mysterious allure and the innovative techniques da Vinci employed to bring the portrait to life.”
43250 ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its timeless beauty and the profound influence it has had on art and culture throughout history.”
Prompt ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle”
2700 ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, King Henry VIII was crowned king of England on January 1st 1485 at Westminster Abbey. King John’s War (1315-14).”
3300 ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, the Normans were able to establish their own rule over England and established a powerful military force that would be able to defeat any other army in Europe at the time.”
5100 ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, King Harold II was crowned as king by his Norman supporters and became known as ’the Conqueror’.”
15700 ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle was an era where the Normans were able to conquer and control much of Europe including France, Spain, Italy, Germany, Poland, Hungary, etc.”
35000 ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, William became King and his descendants would rule over Britain for many years until Henry VIII broke away from Rome’s authority by declaring himself Head of the Church (the Pope).”
43250 ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, William I became King and he was known as William The Conqueror (1066-1087).”
Table II summarizes the averaged scores across the evaluated criteria for each checkpoint. Initially, the model receives middling marks on attributes such as Coherence and Fluency, reflecting a nascent grasp of narrative construction and linguistic form. Over time, as training steps increase, the gradual upward trend in categories like Depth and Completeness suggests that the model becomes not only more reliable at presenting known information but also more adept at elaborating thoughtfully on prompts. Nevertheless, the Factual Accuracy (FA) score remains more modest, with even the most extensively trained checkpoint (43250) receiving only a 5 in this category. This relatively muted improvement in factual accuracy underscores the inherent difficulty of producing historically and contextually precise content, even at later stages of model refinement. It indicates that the model is still not sufficiently trained in factual domains, reinforcing observations that large language models often require extensive training—and targeted data coverage—to reduce hallucinations. This also helps explain why earlier-generation models (such as ChatGPT-3.5 or LLaMa 1 [3] and 2 [10]), trained with fewer tokens, were more prone to factual distortions, whereas newer models (ChatGPT-4 [11], LLaMa 3 [7] and Qwen 2.5 [12]) with more comprehensive training corpora tend to display far fewer factual errors. In essence, these aggregated scores provide a quantitative reflection of a well-known progression in language model development: as training deepens and data coverage broadens, models transition from producing disjointed or factually dubious content toward more coherent, contextually appropriate, and accurate outputs—though even the most advanced checkpoints still leave room for improvement in factual domains.
Table III presents completions for three prompt examples. The first prompt, ”The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note”, is relatively straightforward and factually well-known, focusing on describing a widely recognized historical landmark. The second prompt, ”The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for”, also deals with a familiar cultural artifact whose key attributes (e.g., enigmatic smile, Renaissance painting techniques) are commonly referenced. In contrast, the third prompt, ”The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle”, involves historical specificity and requires more precise factual recall.

### v3

@@ Abstract
Pre-training large language models is a complex endeavor influenced by multiple factors, including model architecture, data quality, training continuity, and hardware constraints. In this paper, we share insights gained from the experience of training DMaS-LLaMa-Lite, a fully open-source, 1.7-billion-parameter, LLaMa-based model, on approximately 20 billion tokens of carefully curated data. We chronicle the full training trajectory, documenting how evolving validation loss levels and downstream benchmarks reflect transitions from incoherent text to fluent, contextually grounded output. Beyond pre-training, we extend our analysis to include a post-training phase focused on instruction tuning, where the model was refined to produce more contextually appropriate, user-aligned responses. We highlight practical considerations such as the importance of restoring optimizer states when resuming from checkpoints, and the impact of hardware changes on training stability and throughput. While qualitative evaluation provides an intuitive understanding of model improvements, our analysis extends to various performance benchmarks, demonstrating how high-quality data and thoughtful scaling enable competitive results with significantly fewer training tokens. By detailing these experiences and offering training logs, checkpoints, and sample outputs, we aim to guide future researchers and practitioners in refining their pre-training strategies. The training script is available on Github and the model checkpoints are available on Huggingface.
keywords:
large language model, natural language processing, deep learning
1 Introduction
Large language models (LLMs) have achieved unprecedented performance across a variety of natural language understanding and generation tasks. These advancements have been driven by innovations in model architectures, pre-training methodologies, and the availability of large-scale, high-quality data (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023a). However, despite significant progress, many aspects of LLM training remain underexplored. Practical challenges—ranging from data curation and training stability to qualitative and quantitative evaluation—play a critical role in determining model performance but are often overlooked in favor of final benchmark results.
In this paper, we present our experience pre-training and post-training DMaS-LLaMa-Lite, a 1.7-billion-parameter LLaMa-based model, on 20 billion tokens of carefully curated data. Unlike many existing studies that focus solely on the final outputs of large-scale training, we emphasize the training process itself and the insights gleaned from it. Specifically, we document the following key aspects of our work:
@@ 1 Introduction
    Post-Training Instruction Tuning: We refine the pretrained model using parameter-efficient fine-tuning, leveraging high-quality instruction datasets. We documented the model’s behavioral change from the process and the model’s sensitivity to the prompt.
Our findings collectively offer insights into the practical challenges and trade-offs encountered during the pre-training of LLMs. By sharing detailed training logs, checkpoint artifacts, and sample completions, we aim to support reproducibility and provide guidance for both researchers and practitioners seeking to optimize their own training strategies. Importantly, this work serves as a foundation for future studies that may uncover additional best practices, methodological refinements, or novel evaluation techniques in the pre-training of large language models. The training script is available on Github [^1] https://github.com/McGill-DMaS/DMaS-LLaMa-Lite-Training-Code. The model checkpoints are available on Huggingface [^2] https://huggingface.co/collections/McGill-DMaS/dmas-llama-lite-6761d97ba903f82341954ceb.
The remainder of this paper is organized as follows: Section  2 describes the model architecture, training setup, and data preprocessing. Section  3 presents our findings from the pre-training experience, including quantitative and qualitative evaluations, and a discussion of broader implications. Section  4 details the post-training phase, including instruction tuning and qualitative analysis. Finally, Section  5 concludes with key takeaways and directions for future work.
2 Model Configuration and Training Setup
2.1 Model Architecture
Our model, DMaS-LLaMa-Lite, is a 1.7B-parameter (1716M) variant of LLaMa (Touvron et al., 2023a) adapted to use a GPT-2 tokenizer. The configuration details are presented in Table 1. The model features 36 transformer layers, each with a hidden size of 2048 and an intermediate size of 5120 in the feed-forward layers. We employ 32 attention heads, grouped into 8 key-value heads, and use a rotary positional embedding scheme. Norm layers use RMS normalization (Zhang and Sennrich, 2019), and the model’s non-linear activation is SiLU (Hendrycks and Gimpel, 2016). The model is trained on tokens produced by the GPT-2 tokenizer.
Table 1: Model configuration for the 1.7B-parameter DMaS-LLaMa-Lite model.
@@ 2.1 Model Architecture
2.2 Training Data and Preprocessing
The training was conducted using a subset of the HuggingFaceFW/fineweb-edu 100BT (Penedo et al., 2024) dataset, derived from the broader FineWeb dataset. FineWeb-Edu comprises approximately 1.3 trillion tokens of high-quality educational content, curated through a meticulous multi-stage filtering process. FineWeb itself is a 15-trillion-token dataset sourced from 96 Common Crawl snapshots, designed to improve LLM performance through advanced filtering and deduplication methods.
FineWeb-Edu emphasizes educational text, selected using a classifier trained on synthetic annotations generated by Llama-3-70B-Instruct (Dubey et al., 2024). This classifier was fine-tuned to prioritize grade-school and middle-school-level knowledge while excluding overly technical content. The dataset employs custom filters targeting document structure and content quality, ensuring a high proportion of semantically coherent and educationally relevant text.
@@ 2.2 Training Data and Preprocessing
2.3 Optimizer and Learning Rate Schedule
@@
2.4 Training Procedure and Checkpoints
The training was conducted on one or two RTX A6000 GPUs, reflecting realistic constraints in research and production environments. When using two GPUs, PyTorch’s DistributedDataParallel (DDP) (Paszke et al., 2019) was employed to ensure efficient data and model parallelism. Transitions in the hardware setup (e.g., moving from a single RTX A6000 GPU to dual RTX A6000 GPUs) served as natural breakpoints, allowing us to examine how changes in training continuity, combined with optimizer state restoration or omission, affected validation loss and model quality.
3 Pre-training
3.1 Validation Loss Trajectories and Training Metrics
3.2 Qualitative Evaluation of Generated Text
Table 2: Generated completions at different training steps.
Checkpoint Generated Text
Prompt "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note"
2700 "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the impressive size and precision of the structures, marveling at how such feats of engineering were accomplished thousands of years ago without modern tools."
3300 "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the mystery surrounding their construction, as well as the stunning alignment with celestial bodies."
5100 "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the grandeur and historical significance of these monuments, which continue to captivate people from around the globe."
15700 "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note their intricate design, which reflects the advanced knowledge of ancient Egyptian engineers and architects."
35000 "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the awe-inspiring scale and the enduring mystery of how they were constructed with such precision."
43250 "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note the breathtaking beauty and the remarkable testament to the ingenuity of ancient civilizations."
Prompt "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for"
2700 "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its enigmatic smile and the use of sfumato technique to create a lifelike expression."
3300 "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its detailed background, which demonstrates da Vinci’s mastery of perspective."
5100 "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its captivating expression, which has intrigued viewers and art historians for centuries."
15700 "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for the subtle play of light and shadow on the subject’s face, a hallmark of Renaissance art."
35000 "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its mysterious allure and the innovative techniques da Vinci employed to bring the portrait to life."
43250 "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for its timeless beauty and the profound influence it has had on art and culture throughout history."
Prompt "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle"
2700 "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, King Henry VIII was crowned king of England on January 1st 1485 at Westminster Abbey. King John’s War (1315-14)."
3300 "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, the Normans were able to establish their own rule over England and established a powerful military force that would be able to defeat any other army in Europe at the time."
5100 "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, King Harold II was crowned as king by his Norman supporters and became known as ’the Conqueror’."
15700 "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle was an era where the Normans were able to conquer and control much of Europe including France, Spain, Italy, Germany, Poland, Hungary, etc."
35000 "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, William became King and his descendants would rule over Britain for many years until Henry VIII broke away from Rome’s authority by declaring himself Head of the Church (the Pope)."
43250 "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle, William I became King and he was known as William The Conqueror (1066-1087)."
Table 3 summarizes the averaged scores across the evaluated criteria for each checkpoint. Initially, the model receives middling marks on attributes such as Coherence and Fluency, reflecting a nascent grasp of narrative construction and linguistic form. Over time, as training steps increase, the gradual upward trend in categories like Depth and Completeness suggests that the model becomes not only more reliable at presenting known information but also more adept at elaborating thoughtfully on prompts. Nevertheless, the Factual Accuracy (FA) score remains more modest, with even the most extensively trained checkpoint (43250) receiving only a 5 in this category. This relatively muted improvement in factual accuracy underscores the inherent difficulty of producing historically and contextually precise content, even at later stages of model refinement. It indicates that the model is still not sufficiently trained in factual domains, reinforcing observations that large language models often require extensive training—and targeted data coverage—to reduce hallucinations. This also helps explain why earlier-generation models (such as ChatGPT-3.5 or LLaMa 1 (Touvron et al., 2023a) and 2 (Touvron et al., 2023b)), trained with fewer tokens, were more prone to factual distortions, whereas newer models (ChatGPT-4 (OpenAI, 2023), LLaMa 3 (Dubey et al., 2024) and Qwen 2.5 (Yang et al., 2024)) with more comprehensive training corpora tend to display far fewer factual errors. In essence, these aggregated scores provide a quantitative reflection of a well-known progression in language model development: as training deepens and data coverage broadens, models transition from producing disjointed or factually dubious content toward more coherent, contextually appropriate, and accurate outputs—though even the most advanced checkpoints still leave room for improvement in factual domains.
Table LABEL:tab:battle_of_hastings presents completions for three prompt examples. The first prompt, "The Pyramids of Giza in Egypt are some of the oldest man-made structures in the world. Visitors often note", is relatively straightforward and factually well-known, focusing on describing a widely recognized historical landmark. The second prompt, "The Mona Lisa, painted by Leonardo da Vinci, is one of the most famous artworks in the world. It is notable for", also deals with a familiar cultural artifact whose key attributes (e.g., enigmatic smile, Renaissance painting techniques) are commonly referenced. In contrast, the third prompt, "The Battle of Hastings, a key event in English history, occurred in the year 1066. It resulted in significant changes to the ruling class of England. Following this battle", involves historical specificity and requires more precise factual recall.
By checkpoint 35000, the text becomes significantly more coherent and contextually relevant, though it introduces minor inaccuracies, such as the association with Henry VIII. At checkpoint 43250, the output demonstrates marked improvement, correctly identifying William I as "William The Conqueror" and providing accurate dates and titles.
These observations align with the scores shown in Table 3: less sufficiently trained language models are more prone to “hallucination”—the production of factually incorrect or contextually inconsistent details—while more extensively trained models show a marked reduction in such errors. As the model’s training proceeds, it appears to acquire a stronger grounding in historical context, produce more thematically coherent narratives, and adhere more closely to factual accuracy. In other words, increased training not only enhances fluency and coherence but also mitigates the tendency to fabricate details, resulting in outputs that are both stylistically refined and factually sound.
In addition to subjective assessment, we further analyzed the relationship between the average qualitative score and three key variables: training steps, validation loss, and Hellaswag (Zellers et al., 2019) scores (see Table 4 for the progression of these metrics). Correlation analysis yielded the following results:
Table 3: Evaluation Scores for Different Training Steps
@@ 3.2 Qualitative Evaluation of Generated Text
3.3 Comparison and Discussion
The performance of DMaS-LLaMa-Lite is compared against TinyLLaMa’s 2T checkpoint (Zhang et al., 2024), a series of lightweight LLaMa-based models trained with significantly larger token counts. TinyLLaMa’s 2T checkpoint, a 1.1B-parameter model trained on 2 trillion tokens using the SlimPajama dataset (Soboleva et al., 2023), represents an established baseline for small-scale LLMs.
The benchmarks used for evaluation are as follows:
  1. 1.
    ARC Challenge: A subset of the AI2 Reasoning Challenge (Clark et al., 2018), which focuses on grade-school science questions that require reasoning and commonsense knowledge.
  2. 2.
  3. 3.
    BoolQ: A yes/no question-answering benchmark (Clark et al., 2019) that requires reading comprehension and contextual reasoning.
  4. 4.
    HellaSwag: A commonsense reasoning benchmark (Zellers et al., 2019) where the goal is to select the most plausible continuation of a given sentence.
  5. 5.
    OpenBookQA: A multiple-choice question benchmark (Mihaylov et al., 2018) that requires reasoning over a small “open book” of facts.
  6. 6.
    PIQA: The Physical Interaction QA benchmark (Bisk et al., 2020), which tests commonsense physical reasoning, such as how objects are used.
  7. 7.
    Winogrande: A large-scale Winograd Schema Challenge (Sakaguchi et al., 2020) dataset, focusing on resolving pronoun ambiguity using commonsense reasoning.